Show that a field f may be considered as a vector space over f if scalar multiplication is identified with field multiplication



1
Expert's answer
2021-12-12T15:36:55-0500

Let V = { (a) | a in F } 


To prove; we show that vector addition is commutative.


Commutativity of addition, Here x, y are in V


(x)+(y) = (y)+(x)


(x+ y)=(y+ x): vector addition

x + (X+ y) = (X + x)+y: associative property


"\therefore" a field f may be considered as a vector space over f